# The Three Regimes of Offline-to-Online Reinforcement Learning

### v1

@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive successes in a variety of domains (Mnih et al., 2015; Silver et al., 2017; Degrave et al., 2022), but its reliance on large amounts of online interaction often makes direct application to real-world problems challenging. To address this challenge, recent research has turned to leveraging pre-collected datasets for offline learning through offline RL (Levine et al., 2020) or imitation learning (Osa et al., 2018). This paradigm reduces the need for costly or unsafe online exploration by providing a strong initial policy trained entirely from offline data. However, policies trained purely offline are often suboptimal and fail to generalize to states outside the dataset’s support, making online fine-tuning essential. Offline-to-online RL addresses this issue by first pretraining an agent on an offline dataset and then fine-tuning it with additional online interactions to further improve performance.
@@
While offline-to-online RL has led to promising results (Nair et al., 2020; Lee et al., 2022), the online RL fine-tuning has shown highly inconsistent empirical behavior. Design choices that work well in one setting can fail completely in another. For example, as shown in Figure 1, on D4RL tasks (Fu et al., 2020) such as antmaze-large-play-v2, Warm-Start RL (WSRL) (Zhou et al., 2024), which relies on the pretrained policy and discards the offline dataset during online fine-tuning, substantially outperforms RL with Prior Data (RLPD) (Ball et al., 2023), which uses only the offline dataset. In contrast, on D4RL tasks such as relocate-binary-v0, the opposite pattern emerges, with RLPD outperforming WSRL by a wide margin.
These seemingly inconsistent outcomes raise one fundamental question:
@@
To answer this question, we propose a stability–plasticity principle for offline-to-online RL that can explain these seemingly inconsistent outcomes, which is inspired by prior research on plasticity and stability in neuroscience (McClelland et al., 1995) and machine learning  (Kirkpatrick et al., 2017; Wolczyk et al., 2024; Dohare et al., 2024). This principle states that effective fine-tuning requires a careful balance between stability and plasticity. Stability refers to the preservation of useful prior knowledge, ensuring that competencies acquired during pretraining are not substantially degraded. Plasticity, in contrast, denotes the capacity of the model to adapt flexibly and efficiently to new data. Furthermore, we identify two distinct forms of stability in offline-to-online RL: stability around the pretrained policy $\pi_{0}$,which emphasizes preserving knowledge explicitly encoded in the policy parameters, and stability around the offline dataset $\mathcal{D}$, which emphasizes retaining knowledge implicitly encoded in offline data. As stability and plasticity are inherently in trade-off, this distinction indicates that fine-tuning is more efficient when stability is enhanced with respect to the stronger source of prior knowledge, whether it be the pretrained policy or the offline dataset.
Building on this insight, we propose a taxonomy of three regimes for offline-to-online RL, each capturing a distinct relationship between the pretrained policy and the offline dataset. This framework provides actionable guidance for practitioners. As shown in Figure 2, the three regimes are defined based on which source of prior knowledge is stronger, either the pretrained policy or the offline dataset. Moreover, different fine-tuning methods can be systematically categorized according to whether they enhance stability around $\pi_{0}$, stability around $\mathcal{D}$, or plasticity, and each category can be positioned within these regimes. By first determining the regime, one can select or design fine-tuning strategies that match its stability-plasticity requirements. This yields two practical benefits: it helps choose the most suitable method for each setting rather than applying a single uniform state-of-the-art algorithm, and it narrows the search space by indicating whether one should focus on leveraging the pretrained policy or on exploiting the offline dataset, thereby reducing unnecessary trial-and-error.
To validate this framework, we conduct a large-scale empirical study that covers 21 dataset-task compositions across four domains in D4RL: Mujoco locomotion, AntMaze, Adroit, and Kitchen. Combined with three representative pretraining algorithms, this results in 63 experimental settings. The observed behaviors align closely with the predictions of our framework, supporting its utility as a principled basis for guiding design choices of fine-tuning in offline-to-online RL.
@@
  • We propose a stability–plasticity principle for offline-to-online RL.
  • We develop a taxonomy of three regimes based on the relative performance of $\pi_{0}$ and $\mathcal{D}$.
@@ 2 Preliminary: Offline-to-online RL
Figure 2: Overview of the three regimes in offline-to-online RL, defined based on the relative performance of the pretrained policy $J(\pi_{0})$ and the offline dataset $J(\pi_{\mathcal{D}})$. For each regime, our framework indicates which property is most needed during fine-tuning. The boxes at the right show representative design choices that implement these enhancing stability or plasticity.
Consider an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where the performance of a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ is measured by its expected discounted return: $J(\pi)=\mathbb{E}_{\pi,\mathcal{M}}[\sum_{t}\gamma^{t}r_{t}]$. Offline-to-online RL begins by pretraining the agent on an offline dataset $\mathcal{D}$, which is collected from $\mathcal{M}$ under an unknown behavior policy (or mixture of policies), using an offline RL algorithm $A_{\text{off}}$. This yields an offline pretrained agent whose policy is given by $\pi_{0}=A_{\text{off}}(\mathcal{D})$, The fine-tuning step consists in using an algorithm $A_{\text{on}}$ using $\pi_{0}$ to interact with $\mathcal{M}$ in order to derive a final policy $\pi_{N}$:
@@
The offline dataset $\mathcal{D}$ may be reused (Nakamoto et al., 2023) or discarded (Zhou et al., 2024) during fine-tuning. The objective of offline-to-online RL is to co-design $(A_{\text{off}},A_{\text{on}})$ such that the final policy $\pi_{N}$ maximizes $J(\pi_{N})$. In this work, we focus on understanding and improving the online RL fine-tuning component $A_{\text{on}}$ by fixing the offline pretraining component $A_{\text{off}}$.
@@ 3 A Stability–Plasticity Principle on Knowledge Acquisition
This section introduces a formal framework for reasoning about fine-tuning in offline-to-online RL. Our goal is to characterize when and how online training leads to improvements over the offline initialization—without degrading what was already learned. We define two complementary properties of online fine-tuning: stability, the ability to preserve previously acquired performance, and plasticity, the capacity to improve further. These are grounded in the notion of a knowledge level, understood as the expected return encoded either in the dataset or in the pretrained policy. We show that final performance admits a decomposition into three terms—prior knowledge, stability, and plasticity. We leverage this perspective provides both theoretical insight and practical guidance.
@@ 3.1 Quantifying Prior Knowledge
We distinguish two sources of prior knowledge available before online fine-tuning. The first is the knowledge encoded in the dataset, captured by the performance of the behavior policy that generated it. The second is the knowledge embodied in the pretrained policy, obtained by running an offline RL algorithm on the dataset. Both can serve as baselines for measuring stability and plasticity.
Knowledge from the dataset: $J(\pi_{\mathcal{D}})$.
Let $\mathcal{D}$ be the offline dataset and $\pi_{\mathcal{D}}$ be an abstract behavior policy representing the data-generating process. While $\mathcal{D}$ may be collected from a mixture of policies, $\pi_{\mathcal{D}}$ serves as a convenient abstraction. Its performance can be estimated by the average return:
$J(\pi_{\mathcal{D}})\approx\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}r_{i,t}.$
This measure provides a scalar summary of the return level encoded in the dataset. It does not capture coverage or diversity, but it defines a baseline that fine-tuning should not fall below.
Knowledge from the pretrained policy: $J(\pi_{0})$.
Let $\pi_{0}$ be the policy produced by applying an offline RL algorithm $A_{\text{off}}$ to $\mathcal{D}$. Its performance, $J(\pi_{0})$, reflects not only the quality of the dataset but also the inductive biases of $A_{\text{off}}$ and the difficulty of the MDP $\mathcal{M}$. This policy can provide a strong initialization for online fine-tuning, though it does not necessarily dominate the dataset baseline. We therefore consider both $J(\pi_{\mathcal{D}})$ and $J(\pi_{0})$ jointly as candidate sources of prior knowledge.
@@ 3.2 Knowledge Decomposition and the Three Regimes of Offline-to-online RL
Stability with regard to a knowledge level.
We define the stability of an online RL training process with respect to a knowledge level $l$, as the ability to retain the relative performance:
$\text{Stability}(l)=\min\left(\min_{0\leq n\leq N}J(\pi_{n})-l,0\right).$ (2)
@@
$J^{*}_{\text{off}}=\max\left(J(\pi_{0}),\;J(\pi_{\mathcal{D}})\right).$
We refer to this as the offline performance baseline, and we define stability with respect to it as:
@@
Plasticity.
We define the plasticity of an online RL algorithm as the ability to acquire new knowledge:
$\text{Plasticity}=\max_{0\leq i\leq N}J(\pi_{i})-\min_{0\leq j\leq N}J(\pi_{j})\geq 0.$ (4)
@@
Our definition enables such comparison by relating both concepts through a knowledge decomposition:{myframe}
@@
$\underbrace{\max_{0\leq n\leq N}J(\pi_{n})}_{\text{Final knowledge}}=\underbrace{J^{*}_{\text{off}}}_{\text{Prior knowledge}}+\underbrace{\text{Stability}(J^{*}_{\text{off}})}_{\text{Degradation on prior knowledge}\leq 0}+\underbrace{\text{Plasticity}}_{\text{Online knowledge}\geq 0}.$
@@
This equation states that the final knowledge an agent acquires after offline pretraining and online fine-tuning is the outcome of three interacting components. The first term is the prior knowledge provided by the offline phase, either through the dataset or the pretrained policy, whichever is stronger. The second term measures stability, which records whether this prior knowledge is preserved or degraded during fine-tuning; it is always non-positive since performance can at best be maintained but not exceeded by this term. The third term captures plasticity, the additional knowledge acquired through online interaction, which is non-negative by definition.
Viewed this way, the outcome of offline-to-online RL is neither purely a continuation of offline learning nor purely the result of online finetuning. It is the byproduct of what was preserved from the offline baseline and what was added through adaptation, minus any performance lost in the process.
The three regimes of offline-to-online RL.
@@
With prior knowledge fixed, the objective of online fine-tuning is to improve final performance by maintaining stability with respect to $\max(J(\pi_{0}),J(\pi_{\mathcal{D}}))$ while ensuring sufficient plasticity. Based on this perspective, we identify three regimes for the online fine-tuning phase of offline-to-online RL: Superior: where $J(\pi_{0})>J(\pi_{\mathcal{D}})$; Comparable: where $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$; Inferior: where $J(\pi_{0})<J(\pi_{\mathcal{D}})$. These regimes are intended to reflect substantial differences in $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$, since small performance gaps may not be meaningful and therefore should not determine regime assignment.
@@
This regime taxonomy provides a principled framework for reasoning about the stability–plasticity trade-off in offline-to-online RL. It clarifies which source of prior knowledge should anchor stability in a given setting, as shown in Figure 2. In the Superior Regime, stability relative to $\pi_{0}$ should be prioritized, meaning that online updates should avoid degrading below the performance already achieved by the pretrained policy. In the Inferior Regime, stability relative to $\mathcal{D}$ should be emphasized, meaning that fine-tuning should at least preserve the return level encoded in the dataset. In the Comparable Regime, both baselines provide similar performance, so preserving either suffices. By categorizing design choices according to the stability or plasticity they promote, our framework turns what previously appeared as a disconnected set of practices into a structured and predictable landscape, providing a principled foundation for future algorithm design.
@@ 4 Design Choices in Improving Stability or Plasticity
Minimal baseline.
We begin with defining a naive online RL fine-tuning baseline that serves as the reference point for introducing additional components, which is intentionally minimalist. It applies a standard online RL algorithm initialized with an offline-pretrained agent, without any further modifications. This baseline anchors the analysis and makes the marginal effect of each added component interpretable.

### v2

@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive successes in a variety of domains (mnih2015human; AlphaGo_Zero; degrave2022magnetic), but its reliance on large amounts of online interaction often makes direct application to real-world problems challenging. To address this challenge, recent research has turned to leveraging pre-collected datasets for offline learning through offline RL (levine2020offline) or imitation learning (osa2018algorithmic). This paradigm reduces the need for costly or unsafe online exploration by providing a strong initial policy trained entirely from offline data. However, policies trained purely offline are often suboptimal and fail to generalize to states outside the dataset’s support, making online fine-tuning essential. Offline-to-online RL addresses this issue by first pretraining an agent on an offline dataset and then fine-tuning it with additional online interactions to further improve performance.
@@
While offline-to-online RL has led to promising results (nair2020awac; lee2022offline), online RL fine-tuning suffers from highly inconsistent empirical behavior: design choices that work well in one setting can fail completely in another. For example, as shown in Figure 1, on D4RL tasks (fu2020d4rl) such as antmaze-large-play-v2, Warm-Start RL (WSRL) (zhou2024efficient), which relies on the pretrained policy and discards the offline dataset during online fine-tuning, substantially outperforms RL with Prior Data (RLPD) (ball2023efficient), which uses only the offline dataset. In contrast, on D4RL tasks such as relocate-binary-v0, the opposite pattern emerges, with RLPD outperforming WSRL by a wide margin. These seemingly inconsistent outcomes raise one fundamental question:
@@
To answer this question, we propose a stability–plasticity principle for offline-to-online RL that can explain these seemingly inconsistent outcomes, which is inspired by prior research on plasticity and stability in neuroscience (mcclelland1995there) and machine learning  (kirkpatrick2017overcoming; wolczyk2024fine; dohare2024loss). Guided by the principle, effective fine-tuning requires a careful balance between stability and plasticity. Stability refers to the preservation of useful prior knowledge, ensuring that competencies acquired during pretraining are not substantially degraded. Plasticity, in contrast, denotes the capacity of the model to adapt flexibly and efficiently to new data. Furthermore, we identify two distinct forms of stability in offline-to-online RL: stability around the pretrained policy $\pi_{0}$, which emphasizes preserving knowledge explicitly encoded in the policy parameters, and stability around the offline dataset $\mathcal{D}$, which emphasizes retaining knowledge implicitly encoded in offline data. As stability and plasticity are inherently in trade-off, this distinction indicates that fine-tuning is more efficient when stability is enhanced with respect to the stronger source of prior knowledge, whether it is the pretrained policy or the offline dataset.
Building on this insight, we propose a taxonomy of three regimes for offline-to-online RL, each capturing a distinct relationship between the pretrained policy and the offline dataset. As shown in Figure 2, the three regimes are defined based on which source of prior knowledge is stronger, either the pretrained policy or the offline dataset. Moreover, different fine-tuning methods can be systematically categorized according to whether they enhance stability around $\pi_{0}$, stability around $\mathcal{D}$, or plasticity. By first determining the regime, one can select or design fine-tuning strategies that match its stability-plasticity requirements. This yields two practical benefits: it helps choose the most suitable method for each setting rather than applying a single uniform state-of-the-art algorithm, and it narrows the search space by indicating whether one should focus on leveraging the pretrained policy or on exploiting the offline dataset, thereby reducing unnecessary trial-and-error.
To validate this framework, we conduct a large-scale empirical study that covers 21 dataset-task compositions across four D4RL domains (MuJoCo locomotion, AntMaze, Adroit, and Kitchen) and three representative pretraining algorithms, yielding 63 settings. The results on stability, plasticity, and final performance, align closely with the predictions of our framework, supporting its utility as a principled basis for guiding design choices of fine-tuning in offline-to-online RL.
@@
  • We propose a stability–plasticity principle for offline-to-online RL that decomposes the fine-tuning performance into quantifiable stability and plasticity.
  • Grounded in this principle, we develop a framework featuring a three-regime taxonomy defined by the relative performance of $\pi_{0}$ and $\mathcal{D}$.
@@ 2 Preliminary: Offline-to-Online RL
Figure 2: Overview of the three regimes in offline-to-online RL, defined based on the relative performance of the pretrained policy $J(\pi_{0})$ and the offline dataset $J(\pi_{\mathcal{D}})$. For each regime, our framework indicates which property is most needed during fine-tuning. The boxes at the right show representative design choices that implement these enhancing stability or plasticity. Dashed arrows denote weaker connections than solid arrows.
Consider an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where the performance of a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ is measured by its expected discounted return: $J(\pi)=\mathbb{E}_{\pi,\mathcal{M}}[\sum_{t}\gamma^{t}r_{t}]$. Offline-to-online RL begins by pretraining the agent on an offline dataset $\mathcal{D}$, which is collected from $\mathcal{M}$ under an unknown behavior policy (or mixture of policies), using an offline RL algorithm $A_{\text{off}}$. This yields an offline pretrained agent whose policy is given by $\pi_{0}=A_{\text{off}}(\mathcal{D})$. The fine-tuning step consists of using an online algorithm $A_{\text{on}}$ that starts from $\pi_{0}$ and interacts with $\mathcal{M}$ to obtain a final policy $\pi_{N}$ after $N$ iterations:
@@
The offline dataset $\mathcal{D}$ may be reused (nakamoto2023cal) or discarded (zhou2024efficient) during fine-tuning. The objective of offline-to-online RL is to co-design $(A_{\text{off}},A_{\text{on}})$ such that the final policy $\pi_{N}$ maximizes $J(\pi_{N})$. In this work, we focus on understanding and improving the online RL fine-tuning component $A_{\text{on}}$ by fixing the offline pretraining component $A_{\text{off}}$.
@@ 3 A Stability–Plasticity Principle on Knowledge Acquisition
This section introduces a formal framework for reasoning about fine-tuning in offline-to-online RL. Our goal is to characterize when and how online training leads to improvements over the offline initialization or degradations of what was already learned. We define two complementary properties of online fine-tuning: stability, the ability to preserve previously acquired performance, and plasticity, the capacity to improve further. These are grounded in the notion of a knowledge level, understood as the expected return encoded either in the dataset or in the pretrained policy. We show that final performance admits a decomposition into three terms — prior knowledge, stability, and plasticity. We leverage this perspective provides both theoretical insight and practical guidance.
@@ 3.1 Quantifying Prior Knowledge
We distinguish two sources of prior knowledge available before online fine-tuning. The first is the knowledge encoded in the dataset, captured by the performance of the behavior policy that generated it. The second is the knowledge embodied in the pretrained policy, obtained by running an offline RL algorithm on the dataset. Both can serve as baselines for measuring stability.
Knowledge from the dataset: $J(\pi_{\mathcal{D}})$. Let $\mathcal{D}$ be the offline dataset and $\pi_{\mathcal{D}}$ be an abstract behavior policy representing the data-generating process. While $\mathcal{D}$ may be collected from a mixture of policies, $\pi_{\mathcal{D}}$ serves as a convenient abstraction. Its performance can be estimated by the average return:
$J(\pi_{\mathcal{D}})\approx\frac{1}{L}\sum_{k=1}^{L}\sum_{t=1}^{T}r_{k,t},$
where $L$ is the number of trajectories in $\mathcal{D}$, and $r_{k,t}$ is the reward at time step $t$ in the $k$-th trajectory. This measure provides a scalar summary of the return level encoded in the dataset. [^1] We do not assume access to additional quantities such as data coverage or performance surrogates (e.g., dense rewards for sparse-reward tasks), which could further enrich the quantification of prior knowledge.
Knowledge from the pretrained policy: $J(\pi_{0})$. Let $\pi_{0}$ be the policy produced by applying an offline RL algorithm $A_{\text{off}}$ to $\mathcal{D}$. Its performance, $J(\pi_{0})$, reflects not only the quality of the dataset but also the inductive biases of $A_{\text{off}}$ and the difficulty of the MDP $\mathcal{M}$. This policy can provide a strong initialization for online fine-tuning, though it does not necessarily dominate the dataset baseline. We therefore consider both $J(\pi_{\mathcal{D}})$ and $J(\pi_{0})$ jointly as candidate sources of prior knowledge.
@@ 3.2 Knowledge Decomposition and the Three Regimes of Offline-to-Online RL
Stability with regard to a knowledge level. We define the stability of an online RL training process with respect to a knowledge level $l$, as the ability to retain the relative performance:
$\vskip-10.00002pt\text{Stability}(l)\vcentcolon=\min\left(\min_{0\leq n\leq N}J(\pi_{n})-l,0\right).$ (2)
@@
$J^{*}_{\text{off}}\vcentcolon=\max\left(J(\pi_{0}),\;J(\pi_{\mathcal{D}})\right).$ (3)
We refer to this as the offline performance baseline, and the stability with respect to it is:
@@
Plasticity. We define the plasticity as the ability to acquire new knowledge in online fine-tuning:
$\text{Plasticity}\vcentcolon=\max_{0\leq i\leq N}J(\pi_{i})-\min_{0\leq j\leq N}J(\pi_{j})\geq 0.$ (5)
@@
By relating these concepts through a knowledge decomposition, we have:{myframe}
@@
$\underbrace{\max_{0\leq n\leq N}J(\pi_{n})}_{\text{Final knowledge}}=\underbrace{J^{*}_{\text{off}}}_{\text{Prior knowledge}}+\underbrace{\text{Stability}(J^{*}_{\text{off}})}_{\text{Degradation on prior knowledge}\leq 0}+\underbrace{\text{Plasticity}}_{\text{Online knowledge}\geq 0}.$ (6)
@@
This equation states that the final knowledge an agent acquires after offline pretraining and online fine-tuning is the outcome of three interacting components. (1) The first term is the prior knowledge provided by the offline phase, either through the dataset or the pretrained policy, whichever is stronger. (2) The second term measures stability, which records whether this prior knowledge is preserved or degraded during fine-tuning; it is always non-positive since performance can at best be maintained but not exceeded by this term. (3) The third term captures plasticity, the additional knowledge acquired through online interaction, which is non-negative by definition. Therefore, the improvement over the prior knowledge is given by the sum of plasticity and stability.
@@
The three regimes of offline-to-online RL. With prior knowledge fixed, the objective of online fine-tuning is to improve final performance by maintaining stability with respect to $\max(J(\pi_{0}),J(\pi_{\mathcal{D}}))$ while ensuring sufficient plasticity. Based on this perspective, we identify three regimes for the online fine-tuning phase of offline-to-online RL: Superior: where $J(\pi_{0})>J(\pi_{\mathcal{D}})$; Comparable: where $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$; Inferior: where $J(\pi_{0})<J(\pi_{\mathcal{D}})$. These regimes are intended to reflect substantial differences in $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$, since small performance gaps may not be meaningful and therefore should not determine regime assignment. We additionally explore alternative taxonomies using different metrics; the full details are given in Appendix C.
@@
This regime taxonomy provides a principled framework for reasoning about the stability–plasticity trade-off in offline-to-online RL. It clarifies which source of prior knowledge should anchor stability in a given setting, as shown in Figure 2. In the Superior Regime, stability relative to $\pi_{0}$ should be prioritized, because $\pi_{0}$ contains more useful knowledge than $\mathcal{D}$. In the Inferior Regime, stability relative to $\mathcal{D}$ should be emphasized, as $\mathcal{D}$ contains more useful knowledge than $\pi_{0}$. In the Comparable Regime, both baselines provide similar knowledge, so preserving either helps. At the same time, maintaining sufficient plasticity across all regimes is essential for efficient fine-tuning. By categorizing design choices according to the stability or plasticity they promote, our framework turns what previously appeared as a disconnected set of practices into a structured and predictable landscape, providing a principled foundation for future algorithm design.
@@ 4 Design Choices in Improving Stability or Plasticity
Minimal baseline. We begin with defining a naive online RL fine-tuning baseline that serves as the reference point for introducing additional components, which is intentionally minimalist. It applies a standard online RL algorithm (e.g., SAC (haarnoja2018soft)) initialized with an offline-pretrained agent, without any further modifications. This baseline anchors the analysis and makes the marginal effect of each added component interpretable.